I've invested $800 for 7 years at 6% simple interest.
A) How much interest was earned?

B) What will the total value of the investment be when it matures at the end of the 7 years?

Answer by Theo(13342)

6% simple interest formula is:

i = p * r * n

i = interest
p = principal
r = interest rate per time period
n = number of time periods.

formula becomes:

i = 800 * .06 * 7 = 336.

value of the investment will be 800 + 336 = 1136.